three brothers have ages that are consecutive odd intergers. the difference between the age of the oldest and twice the age of the youngest is twenty four less than the middle brothers age. find their ages
Answer by vicgonzerx(31)

X=YOUNGEST
X+2=MIDDLE
X+4=OLDEST
X+4-2X=X+2-24
2X=26
X=13
X+2=15
X+4=17
THE AGE ARE 13,15,17